OpenStudy (anonymous):

A ball is thrown upward with a speed of 40 feet per second from the edge of a cliff 500 feet above the ground. What is the speed of the ball when it hits the ground? Use acceleration due to gravity as –32 feet per second squared and approximate your answer to 3 decimal places.

OpenStudy (anonymous):

@zepdrix @IrishBoy123 @SithsAndGiggles

OpenStudy (anonymous):

a(t)=-32 v(t)=-32t+C v(0)=-32(0)+C 40=C v(t)=-32t+40 What now?

zepdrix (zepdrix):

Anti-Differentiate again to get your position function I suppose :)

terenzreignz (terenzreignz):

Did you miss me? :D Integrate it again. To get the position function, y(t).

OpenStudy (anonymous):

Of course I did :p Are we not looking for speed? Why would I need position, I did that at first and got an insanely high number.

zepdrix (zepdrix):

And note that they gave you another piece of initial data for the new unknown constant that shows up: s(0)=500

zepdrix (zepdrix):

You need to know at what time the object hits the ground, then you can use that time to calculate velocity

OpenStudy (anonymous):

I now have -16t^2+40t+00

terenzreignz (terenzreignz):

@zepdrix so fast T.T

OpenStudy (irishboy123):

quickest, use \(v^2 = u^2 + 2ax\)

zepdrix (zepdrix):

So you're looking for the time t when your object is at a height of 0. That is when it hits the ground. 0=-16t^2+40t+500

OpenStudy (anonymous):

I meant 500 not 00 btw

zepdrix (zepdrix):

These guys and their fancy physics shortcuts :) lol

OpenStudy (anonymous):

Eh I prefer calc over physics

OpenStudy (anonymous):

Leme finish it up.

OpenStudy (anonymous):

\[v^2-u^2=2gh,0^2-(40)^2=2*-32*h,h=\frac{ -40*40 }{ -2*32 }=25\] so it goes 25 feet higher. Now total distance =500+25=525 ft initial speed u=o ft/s using above formula again you can find v ,the speed when it touches the ground.

OpenStudy (anonymous):

It is position. 500ft cliff above the ground=+500 Thrown upward with speed of 40 feet per second-it will eventually fall down to the ground which is negative=40t acceleration is negative because velocity is negative -32 feet per second squared t is seconds so second squared=-32t^2

OpenStudy (anonymous):

Position s=0 means it is a at ground. You find t to know the time it took to reach the ground. You then used the velocity function by taking derivative of position. Plug that time it reached the ground to the velocity function.

OpenStudy (anonymous):

∫v(t)=s(t)=−16t2+40t+500

OpenStudy (anonymous):

when it comes down g is positive or g=32 ft/s^2

OpenStudy (anonymous):

I got t=6.978

OpenStudy (anonymous):

Since other solution is negative

OpenStudy (irishboy123):

well: \(\ddot x = a, \implies \dot x \dfrac{d \dot x}{dx} = a \implies \left|\dfrac{\dot x^2}{2} \right |_u^v = ax \) to meet all those calculus cravings!!

OpenStudy (anonymous):

It is canceled correct?

OpenStudy (anonymous):

Yall spewing out a bit too much enough

zepdrix (zepdrix):

t=7? good

OpenStudy (anonymous):

Is that the final answer or would I simply take t at 0 and plug it into v(t)?

OpenStudy (anonymous):

it should be more than 40

OpenStudy (anonymous):

Which was 7

zepdrix (zepdrix):

correct, plug it in, v(7)=?

OpenStudy (anonymous):

I am getting -184

OpenStudy (anonymous):

It would be 184 ft/sec correct?

zepdrix (zepdrix):

184 for `speed` -184 is the velocity though the negative tells us direction, that it's going down

zepdrix (zepdrix):

I think the original question asked for velocity, ya?

zepdrix (zepdrix):

Oh it said speed :) my bad

OpenStudy (anonymous):

Nope speed

OpenStudy (anonymous):

:)

zepdrix (zepdrix):

Oh, you've got your units correct also, cool!

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

Thanks all

OpenStudy (anonymous):

\[v^2-0^2=2*32*525,v=\sqrt{64*525}=8*5\sqrt{21}=40\sqrt{21}\]

OpenStudy (irishboy123):

@surjithayer +1 !!